OpenStudy (anonymous):

A juice container is a square prism with base edge length 4 inches. When an 8-inch straw is inserted into the container as shown, exactly 1 inch remains outside the container. The length of AB is ___inches, to the nearest inch.

OpenStudy (abmon98):

Could you draw the shape of that square prism

OpenStudy (abmon98):

i think we would use pythagros theorm method

OpenStudy (radar):

Follow the suggestion of @Abmon98. first compute the length of the diagonal for the 4 X 4 square bottom. CB = the square root of the sum of the two sides squared. \[CB=\sqrt{4^{2}+4^{2}}=\sqrt{32}=4\sqrt{2}\] Now solve wait, wait.....if AB is the length of the straw inside the container, it is 8-1 or 7 inches. I thought they were asking for the height of the container.

OpenStudy (radar):

if it was the height , it would be the\[\sqrt{49-32 }=\sqrt{17}\]For some reason that seems too short, but I don't see any errors.